OpenStudy (anonymous):

Is the expression x^3*x^3*x^3 equivalent to x^3*3*3? Why or why not? Explain your reasoning.

OpenStudy (drigobri):

Is the expression x³ * x³ * x³ is equivalent to x³*³*³, because you have a common denominator, so you can work just with the exponents. Or writing in an easier way x³ = x * x * x if you multiply x³ * x³, you will: x³ * x³ (x * x * x) * (x * x * x) = (x * x * x * x * x * x * x * x * x) x^9, right? Them x^9 * x³ (x * x * x * x * x * x * x * x * x) * (x * x * x) = x^27 = x(3*3*3)